# GSRender: Deduplicated Occupancy Prediction via Weakly Supervised 3D Gaussian Splatting

### v1

@@ Abstract
3D occupancy perception is gaining increasing attention due to its capability to offer detailed and precise environment representations. Previous weakly-supervised NeRF methods balance efficiency and accuracy, with mIoU varying by 5-10 points due to sampling count along camera rays. Recently, real-time Gaussian splatting has gained widespread popularity in 3D reconstruction, and the occupancy prediction task can also be viewed as a reconstruction task. Consequently, we propose GSRender, which naturally employs 3D Gaussian Splatting for occupancy prediction, simplifying the sampling process. In addition, the limitations of 2D supervision result in duplicate predictions along the same camera ray. We implemented the Ray Compensation (RC) module, which mitigates this issue by compensating for features from adjacent frames. Finally, we redesigned the loss to eliminate the impact of dynamic objects from adjacent frames. Extensive experiments demonstrate that our approach achieves SOTA(state-of-the-art) results in RayIoU (+6.0), while narrowing the gap with 3D supervision methods. Code will be released soon.
1 Introduction
Figure 1: The limitations of RenderOcc[35]. (a) NeRF-based method face a trade-off between efficiency and performance. (b) Duplicate predictions caused by the uncertainty in depth estimation.
@@ 1 Introduction
Occupancy perception[48], a key focus within the rapidly advancing field of autonomous driving perception, has garnered significant attention from both industry and academia. This task models the entire 3D scene as a collection of grid cells, each assigned relevant attributes such as semantics and motion flow. Compared to traditional 3D object detection[37, 54, 20] and BEV(Bird’s Eye View) perception[32, 12], 3D occupancy perception provides a more fine-grained understanding of the environment and compensates for the inability of BEV space to capture height information. Researchers in this field are actively working to develop solutions that are both efficient and capable of real-time processing.
Most existing approaches[28, 21, 16, 14, 51, 26, 22, 41] heavily rely on BEV perception with their performance closely tied to the accuracy of 3D labels. However, as noted in RenderOcc[35], purifying around 30,000 frames of 3D occupancy labels requires about 4,000 human hours. Additionally, inconsistencies in 3D label generation across different benchmarks further hinder the practical application of 3D occupancy prediction in real-world scenarios, leading to unpredictable model performance across various environments. Consequently, the question of how to diminish or even eradicate the model’s reliance on 3D ground truth data stands as a pivotal and urgent challenge that demands our profound contemplation, especially considering the growing demand for advanced 3D perception in technologies like autonomous driving.
RenderOcc[35] represents a pioneering approach by leveraging solely 2D labels to supervise occupancy prediction tasks, effectively bypassing the limitations associated with 3D labeling constraints. Even though the method is relatively superior, there are still two issues that require our attention, as shown in Figure 1. (a) Sampling Trade-off: NeRF-based methods require sampling 3D points along camera rays, but balancing this is tricky. Oversampling creates computational strain, while undersampling can miss key features. This intricate trade-off could result in a fluctuation of accuracy within the range of 5 to 10 percentage points. (b) Duplicated Predictions: Relying on 2D data for supervision often leads to inaccurate depth predictions and duplicate predictions along camera rays. Such duplications hacks the mIoU metric and thereby diminish the practical utility of 3D occupancy prediction when applied to downstream planning and control tasks.
@@
To address these challenges, we introduce GSRender, an end-to-end network that leverages 3D Gaussian Splatting for rendering [19]. 3D Gaussian Splatting enables real-time, high-quality rendering of photorealistic scenes from a set of images, bypassing the need for complex and uncertain sampling processes. The comparison results, as shown in Figure LABEL:fig:1, demonstrate the effectiveness of our approach. Specifically, we incorporate auxiliary supervision from adjacent frames to address the issue of duplicate predictions. Furthermore, to minimize the potential negative impact of dynamic objects on the rendered scenes, we apply differentiated loss weights for static and dynamic objects, achieve significant performance improvements.
@@
We conduct experiments using renowned datasets Nuscenes[3]. Beyond the conventional mIoU metric[30], we also employed RayIoU[42], a novel metric designed to mitigate the influence of thick surfaces. Impressively, our method achieved SOTA results in RayIoU under the category of 2D weakly-supervised methods. Specifically, we summarize our contributions as follows:
@@
    GSRender is the first to conduct a detailed analysis 3D Gaussian Splatting for weakly supervised 3D occupancy prediction, enhancing spatial occupancy accuracy in dynamic environments without complex sampling.
@@
    We introduce the Ray Compensation module which is easily integrable to effectively mitigate duplicate predictions along identical rays and develop a differentiated loss function tailored for dynamic and static objects in the auxiliary frames.
@@
    We achieved SOTA performance in RayIoU metrics among all 2D weakly supervised methods and significantly reduced the performance gap compared to methods utilizing 3D supervision.
2 Related work
2.1 Duplicate Prediction
Monocular depth estimation[2, 34] is fundamental to all tasks in 3D perception, as inaccuracies can critically undermine the overall accuracy of these systems. The challenge of deriving depth from images in vision-centric 3D perception often results in duplicate predictions and false positive detections along camera rays. Although recent studies have resolved this issue within the domain of 3D object detection[24, 6, 27, 15], the problem of duplicate predictions still persists in more fine-grained tasks, particularly in 3D occupancy prediction with 2D weakly supervision. Specifically, features along the same ray are typically supervised by a single pixel, which results in all voxels along the ray being assigned the same semantic label. Our approach tackles occlusions in the current frame by integrating perspectives from adjacent frames, achieving superior performance over RenderOcc with just one additional frame compared to their seven.
2.2 Occupancy prediction
Since 2022 Tesla AI Day[1], the field of occupancy prediction has seen heightened interest. Mainstream perception models predominantly rely on BEV (Bird’s Eye View) perception, they generally adhere to two main paradigms. The first paradigm employs the LSS (Lift-Splat-Shoot) method[36, 21, 51, 22, 15], which predicts depth and unifies features within the ego-vehicle’s coordinate system. Similarly, the second Transformer-based[44] paradigm [23, 25, 53, 46] utilizes deformable cross-attention [55] to incrementally enhance BEV grid features derived from 2D image features. Existing methods primarily utilize dense features, even though most of the actual space is free. Recent innovative methods have effectively sparsified the feature space with notable results[42, 39, 18, 45, 41]. Moreover, employing 2D supervision for 3D occupancy prediction has gained popularity. RenderOcc[35], for example, uses a NeRF-based approach to bridge semantic and density fields with 2D ground truth. OccNeRF[52] and SelfOcc[17] leverage self-supervision to generate 2D ground truth.
2.3 3D Gaussian Splatting
3D Gaussian Splatting[19] has been widely explored and applied in various fields[40, 5, 49, 50, 38, 11, 31, 47], offering real-time capabilities for large-scale scenes reconstruction. It significantly outperforms NeRF in terms of real-time performance and rendering quality, representing a substantial advancement. Essentially, 3D occupancy prediction is a reconstruction task where vision-centric images serve as inputs and the outputs are coarse-grained models of the scene. For such tasks, adopting a reconstruction-based approach may provide a more intuitive and effective solution. Recent efforts have increasingly treated large-scale outdoor scenes as collections of Gaussians. For instance, GaussianFormer[18] enhances each Gaussian’s properties using a transformer combined with the Gaussian-to-Voxel. GaussianBEV[4] replaces the LSS method with Gaussian Splatting to acquire BEV features, while GaussianOcc[10] implements a two-stage approach that integrates cross-view information from adjacent views to minimize sensitivity to camera poses. In our research, we consider each voxel as a Gaussian and similarly project the 3D Gaussian onto 2D for extensive scene reconstruction.
@@ 2.2 Occupancy prediction
3 Method
3.1 Preliminary
3D Gaussian Splatting[19] is a highly popular method in the field of 3D reconstruction, used for real-time reconstruction of both single objects and large-scale scenes. It conceptualizes the entire scene as a collection of Gaussian. Each Gaussian encompasses attributes such as the mean and covariance of the Gaussian distribution, along with opacity and color. The Gaussian probability for each spatial point is calculated as follows:
@@ 3.1 Preliminary
For independent optimization, store the scale as a 3D vector $\textit{s}\in\mathbb{R}^{3}$ and the rotation as a quaternion $\textit{q}\in\mathbb{R}^{4}$. Then, the 3D Gaussians are projected onto the 2D imaging plane using transformation matrices $W$ and Jacobian matrices $J$. The covariance matrix of the 2D Gaussian is computed as follows:
@@
where $\alpha_{k}$ represents the opacity of the $k$th Gaussian, $G_{k}^{2D}(x)$ is the Gaussian probability at pixel position $x$, and $c_{k}$ denotes the color of the $k$-th Gaussian, represented by spherical harmonic coefficients[8]. $K$ denotes the number of Gaussians contributing to pixel $x$.
Problem Defination. we aim to construct a comprehensive 3D occupancy representation of the scene using images captured from surrounding viewpoints. Specifically, for the vehicle at timestamp t, we take N surrounding images $I=\{I_{1},I_{2},\ldots,I_{N}\}$ as input and predict the 3D occupancy grid $V\in\mathbb{R}^{H_{V}\times W_{V}\times Z_{V}\times L}$, where $H_{V}$, $W_{V}$, and $Z_{V}$ denote the resolution of the voxel space, and $L$ represents the number of categories, including free category for unoccupied space. The formulation for this 3D prediction task is articulated as follows:
@@
Specifically, $\mathbb{H}$ generates an occupancy feature $P\in\mathbb{R}^{H_{V}\times W_{V}\times Z_{V}\times C}$, where C denote the feature dimension, capturing spatial structure and temporal relationships within the scene. Following this, the function $\mathbb{F}$ as a head that takes the scene feature representation $P$ as inputs and outputs the final prediction $V$. In GSRender We replace the nerfhead in RenderOcc with gaussianhead, implementing a more efficient method.
3.2 Overall Framework
Figure 2 illustrates the overall framework of our network architecture. We conceptualize the entire scene as an assembly of numerous Gaussian. In the feature extraction stage, to ensure a fair comparison with RenderOcc, we adopt the same structure. In practical tasks, the feature extraction network architecture can be switched freely to other backbones. In Sec. 3.3, we treat each voxel center as the center of a Gaussian and predict the offset of their properties within the Gaussian. Next, we render the semantics logits and depth using 3D Gaussian splatting and optimize the network with 2D supervision. Finally, in Sec. 3.4, we introduce a multi-frame calibration module aimed at reducing false positive detections caused by duplicate predictions.
@@ 3.2 Overall Framework
3.3 Gaussian Render Module
Gaussian Properties Field. We treat each voxel center as the initial point for a Gaussian distribution, as shown in Figure 3, For a given occupancy feature $V$, In addition to predicting the semantic and opacity attributes, we initialize the Gaussian’s mean and scale, and predict offsets for each properties, excluding rotation. In structured voxel spaces, we contend that rotating Gaussian is unnecessary. Instead, adjusting scale suffices to meet the spatial requirements of the entire voxel space. The regular arrangement of voxels in a uniformly distributed and structured environment already provides sufficient geometric and spatial information, rendering rotational adjustments redundant. Our ablation studies have confirmed this approach. The resulting Gaussian properties $G\in\mathbb{R}^{N\times(7+(L-1))}$, where N denotes the total number of Gaussian. For each Gaussian, we have
@@ 3.3 Gaussian Render Module
During the testing phase, we filter out Gaussian with opacity greater as our predicted results. Specifically
$V(\mu)=\begin{cases}\arg\max(C(\mu))&\text{if }o(\mu)\geq\tau\\ \text{empty label}&\text{if }o(\mu)<\tau,\end{cases}$ (7)
where $\tau$ is the opacity threshold for determining whether a Gaussian is considered empty. Since empty voxels are filtered out based on threshold, the actual number of predicted semantic categories is $L-1$, excluding the ’free’ category. Visualization analysis shows that the majority of effective Gaussians closely align with their voxel centers. Thus, we apply the Gaussian semantics directly to the voxels.
Gaussian Rasterizer, in contrast to NeRF[33], obviates the labor-intensive sampling process. It directly renders the Gaussian to the image. We interpret the semantic logits associated with each Gaussian as colors, and apply alpha-blending to render the semantic logits and depth for each pixel. This approach culminates in the generation of a semantic map $S$ and a depth map $D$.
@@
3.4 Ray Compensation Module
RenderOcc[35] introduces Auxiliary Rays to maintain spatial consistency across various viewpoints. Nevertheless, it fails to resolve the duplicated predictions along these rays. We introduce the Ray Compensation module, which provides multiple perspectives for each object at the feature level, significantly reducing the problem of repeated predictions. Specifically as shown in Figure 4. The upper part of figure illustrates the principle of our method. For the current frame $t$, trees that are obscured might be revealed in adjacent frames from the different viewpoints. To obtain features from adjacent time steps, we first assume that the current Occupancy Feature is denoted as $P_{\text{curr}}\in\mathbb{R}^{H_{V}\times W_{V}\times D_{V}\times C}$. We treat the centers of voxels as homogeneous coordinates $O^{\text{curr\_ego}}_{\text{curr}}\in\mathbb{R}^{H_{V}\times W_{V}\times D_{V}\times 4}$, representing the voxel centers in the ego-car coordinate system at the current time. The position of the adjacent time step in the current time step’s coordinate system is computed as follows:
@@ 3.4 Ray Compensation Module
Dynamic Object. It is noteworthy that this approach may still encounter issues, as illustrated in the lower part of Figure4. For trees obscured in the current frame, they may still be blocked by other moving objects in adjacent frames. In extreme cases, these obstructions might occur at every time step. Therefore, we need to minimize the influence of dynamic objects in adjacent frames. This is achieved by applying a weighting factor, alpha, when calculating the loss for adjacent frames, as detailed in the following subsection.
3.5 Loss Function
To maintain consistency with RenderOcc, We also optimize the Gaussian properties using a class-balanced cross-entropy loss $\mathcal{L}^{\text{seg}}$ for occupancy prediction and the SILog loss $\mathcal{L}^{\text{depth}}$ [7] for depth estimation. For the current frame, we compute the loss similarly to RenderOcc as follows:
$\mathcal{L}_{curr}=\mathcal{L}^{seg}_{curr}+\mathcal{L}^{depth}_{curr},$ (10)

### v2

# GSRender: Deduplicated Occupancy Estimation via Weakly Supervised 3D Gaussian Splatting

@@ Abstract
Weakly-supervised 3D occupancy perception is crucial for vision-based autonomous driving in outdoor environments. Previous methods based on NeRF often face a challenge in balancing the number of samples used. Too many samples can decrease efficiency, while too few can compromise accuracy, leading to variations in the mean Intersection over Union (mIoU) by 5-10 points. Furthermore, even with surrounding-view image inputs, only a single image is rendered from each viewpoint at any given moment. This limitation leads to duplicated predictions, which significantly impacts the practicality of the approach. However, this issue has largely been overlooked in existing research. To address this, we propose GSRender, which uses 3D Gaussian Splatting for weakly-supervised occupancy estimation, simplifying the sampling process. Additionally, we introduce the Ray Compensation module, which reduces duplicated predictions by compensating for features from adjacent frames. Finally, we redesign the dynamic loss to remove the influence of dynamic objects from adjacent frames. Extensive experiments show that our approach achieves SOTA results in RayIoU (+6.0), while also narrowing the gap with 3D-supervised methods. This work lays a solid foundation for weakly-supervised occupancy perception. The code is available at https://github.com/Jasper-sudo-Sun/GSRender.
I INTRODUCTION
@@
Autonomous driving has advanced rapidly in recent years. As a key task, occupancy estimation has attracted attention from both academia and industry. This task models the scene as a grid-based collection, where each grid is assigned relevant attributes such as semantics and motion flow. Compared to traditional 3D object detection [qian20223d, zhou2023fastpillars, kim2021survey] and BEV(Bird’s Eye View) perception [ma2024vision, he2024vision], 3D occupancy estimation provides a more fine-grained understanding of the environment and compensates for the inability of BEV space to capture height information. Researchers are committed to advancing the accuracy and real-time efficiency of comprehensive scene estimation, aiming to push the boundaries of understanding and interaction with dynamic environments.
Most existing approaches [liu2023sparsebev, li2023bevstereo, huang2023tri, huang2022bevdet4d, yu2024panoptic, li2023fb, li2023bevdepth, Tang_2024_CVPR] are heavily dependent on 3D ground truth. As noted in RenderOcc [pan2024renderocc], generating 30,000 frames of 3D labels requires approximately 4,000 human hours. Moreover, inconsistencies in 3D label generation across benchmarks further complicate real-world applications. Thus, reducing or eliminating reliance on 3D labels is a critical challenge, particularly as the demand for 3D perception in autonomous driving grows.
Figure 1: The limitations of RenderOcc [pan2024renderocc]. (a) NeRF-based method face a trade-off between efficiency and precision. (b) Duplicate predictions caused by the uncertainty in depth estimation.
RenderOcc [pan2024renderocc] introduces a method that uses LiDAR point clouds to directly supervise predictions, removing the need for post-processing. Even though RenderOcc is relatively superior, there are still two issues that require our attention, as shown in Figure 1. (a) Sampling Trade-off: NeRF-based methods necessitate sampling 3D points along camera rays, but achieving a balanced sampling remains a significant challenge. This complex trade-off may cause accuracy fluctuations ranging from 5 to 10 percentage points. If sampling is insufficient, some positions might be completely undetectable. (b) Duplicated Predictions: Rendering merely relying on single-view images often leads to duplicated predictions along camera rays. Such duplications artificially inflate the mean Intersection over Union (mIoU) metric [tang2024sparseocc], thereby reducing the practical utility of 3D occupancy estimation when applied to downstream tasks.
@@
To address these challenges, we introduce GSRender, an end-to-end framework that leverages 3D Gaussian Splatting [kerbl3Dgaussians], which enables high-quality rendering of photorealistic scenes by utilizing multiple viewpoint images of the same object. But in occupancy estimation tasks, there are no multi-view images; instead, only surrounding-view images are used, with minimal overlap between different viewpoints. Specifically, we achieve multi-view supervision of the same object by incorporating adjacent frames to address the issue of duplicated predictions. Furthermore, to minimize the potential negative impact of dynamic objects on the rendered scenes, we redesign a differentiated loss for static and dynamic objects, which leads to further performance improvements.
@@
We conducted all experiments on the nuScenes [caesar2020nuscenes] dataset. It outperforms RenderOcc in the conventional mIoU [long2015fully] metric. We also employed RayIoU [tang2024sparseocc], a novel metric specifically designed to more accurately measure the model’s true predictive performance. Impressively, our method achieved SOTA results in RayIoU under the category of weakly-supervised methods. Specifically, we summarize our contributions as follows:
@@
    We present GSRender and a detailed explanation of the effectiveness of 3D Gaussian Splatting in the analysis section for the weakly-supervised occupancy estimation task.
@@
    We propose the Ray Compensation module and a refined loss to mitigate duplicated predictions and handle dynamic and static objects in auxiliary frames, enhancing the practicality of weakly supervised methods.
@@
    We achieved SOTA performance in RayIoU metrics among all 2D weakly-supervised methods and substantially reduced the performance gap compared to methods utilizing 3D supervision.
II Related Work
II-1 Voxel-based Scene Representation.
Efficient representations have become a key focus. Voxel-based representations enable fine-grained representations, driving success in LiDAR segmentation [liong2020amvnet, tang2020searching, cheng20212, Tang_2024_CVPR, ye2021drinet++] and 3D scene completion [cao2022monoscene, roldao2020lmscnet, chen20203d]. Since Tesla AI Day 2022 [tesla2022], occupancy perception has gained increasing attention. Mainstream models fall into two paradigms: the first uses the LSS method [philion2020lift, li2023bevstereo, yu2024panoptic, li2023bevdepth, huang2021bevdet], which predicts depth within the ego-vehicle’s coordinate system, while the second employs Transformer-based methods [vaswani2017attention] [li2023voxformer, li2022bevformer, zhang2023occformer, wei2023surroundocc] with deformable attention [zhu2020deformable] to refine BEV features. However, these methods often rely on dense features, whereas most voxels in the space are empty, leading to computational redundancy in real-world scenarios. Recent works have focused on simplifying the feature space, achieving notable results [tang2024sparseocc, shi2024occupancy, huang2024gaussian, wang2024opus, Tang_2024_CVPR].
II-2 Gaussian Scene Representation.
3D Gaussian Splatting [kerbl3Dgaussians] has been applied widely [szymanowicz2024splatter, charatan2024pixelsplat, xu2024grm, yan2024street, shen2024gamba, gao2024gaussianflow, lu20243d, wu20244d], offering real-time scene reconstruction. It outperforms NeRF in both performance and rendering quality. Recent work treats large-scale outdoor scenes as Gaussian collections, including GaussianFormer [huang2024gaussian], which enhances each Gaussian using a transformer, and GaussianBEV [chabot2024gaussianbev], which replaces LSS with Gaussian Splatting for BEV feature extraction. GaussianOcc [gan2024gaussianocc] integrates cross-view information to minimize sensitivity to camera poses. Our research also models outdoor scenes as Gaussians for detailed reconstruction.
II-3 Weakly-Supervised Estimation.
Due to challenges in obtaining 3D labels, many works focus on weakly-supervised paradigms for semantic scene completion. [Wimbauer_2023_CVPR] achieves semantic completion from a single image, and S4C [hayler2024s4c] improves completion by aligning objects across multiple frames. For occupancy estimation, several methods aim to reduce reliance on 3D labels. RenderOcc [pan2024renderocc] uses a NeRF-based approach to bridge semantic and density fields with depth maps from LiDAR. OccNeRF [zhang2023occnerf] and SelfOcc [huang2024selfocc] use self-supervision for generating 2D semantic and depth maps. However, the issue of duplicated predictions from a single viewpoint persists due to the limited overlap between different viewpoints at the same time. We address this issue by introducing temporal supervision, which helps mitigate the problem of duplicated predictions from a single viewpoint.
@@ II Related Work
III Preliminary
We first briefly introduce the essential knowledge of 3D Gaussian Splatting, the problem definition, and the evaluation metrics we use.
III-1 3D Gaussian Splatting.
3D Gaussian Splatting [kerbl3Dgaussians] is a highly popular method in the field of 3D reconstruction, used for real-time reconstruction of both single objects and large-scale scenes. It conceptualizes the entire scene as a collection of Gaussians. Each Gaussian encompasses attributes such as the mean and covariance of the Gaussian distribution, along with opacity and color. The Gaussian probability for each spatial point is calculated as follows:
@@ III Preliminary
For independent optimization, we represent the scale as a 3D vector $\textit{s}\in\mathbb{R}^{3}$ and the rotation as a quaternion $\textit{q}\in\mathbb{R}^{4}$. Then, the 3D Gaussians are projected onto the 2D imaging plane using transformation matrices $W$ and Jacobian matrices $J$. The covariance matrix of the 2D Gaussian is computed as follows:
@@
where $\alpha_{k}$ represents the opacity of the $k$th Gaussian, $G_{k}^{2D}(x)$ is the Gaussian probability at pixel position $x$, and $c_{k}$ denotes the color of the $k$-th Gaussian, represented by spherical harmonic coefficients [fridovich2022plenoxels]. $K$ denotes the number of Gaussians contributing to pixel $x$.
III-2 Problem Definition.
we aim to construct a comprehensive 3D occupancy representation of the scene using images captured from surrounding viewpoints. Specifically, for the vehicle at timestamp $t$, we take $N$ surrounding images $I=\{I_{1},I_{2},\ldots,I_{N}\}$ as input and predict the 3D occupancy grid $V\in\mathbb{R}^{H_{V}\times W_{V}\times Z_{V}\times L}$, where $H_{V}$, $W_{V}$, and $Z_{V}$ denote the resolution of the voxel space, and $L$ represents the number of categories, including free category for unoccupied space. The formulation for this 3D prediction task is articulated as follows:
@@
Specifically, $\mathbb{H}$ generates an occupancy feature $P\in\mathbb{R}^{H_{V}\times W_{V}\times Z_{V}\times C}$, where $C$ denote the feature dimension, capturing spatial structure and temporal relationships within the scene. Following this, the function $\mathbb{F}$ serves as a head that takes the scene feature representation $P$ as input and outputs the final prediction $V$. RenderOcc [pan2024renderocc] employs a NeRF-based rendering head and achieves multi-view alignment through Auxiliary Rays. In GSRender, we replace the head with a Gaussian-based head and refine the auxiliary rays module, implementing a more efficient approach.
@@
III-3 Evaluation metrics.
We utilized both the mIoU metric and the newly proposed RayIoU [tang2024sparseocc] metric to evaluate our method. The RayIoU metric simulates LiDAR by projecting query rays into the predicted 3D occupancy. A query ray is considered a true positive (TP) if its class matches and the $L1$ depth error between the ground-truth and predicted depths is below a threshold (e.g., 2m). The query ray can originate from the LiDAR position at the current, past, or future moments along the ego path. Temporal casting allows us to evaluate scene completion performance while maintaining a well-posed task. Given C classes, RayIoU is computed as follows:
$\text{RayIoU}=\frac{1}{C}\displaystyle\sum_{c=1}^{C}\frac{\text{TP}_{c}}{\text{TP}_{c}+\text{FP}_{c}+\text{FN}_{c}},$ (6)
where $\text{TP}_{c}$, $\text{FP}_{c}$, and $\text{FN}_{c}$ correspond to the number of true positive, false positive, and false negative predictions for class $c_{i}$. We believe this evaluation method better reflects the model’s true performance, so we focus on using it as our primary evaluation metric.
IV Method
IV-A Overall Framework.
Figure 2 illustrates our overall framework. We conceptualize the entire scene as an assembly of numerous Gaussian. In the feature extraction stage, we adopt the same backbone to ensure a fair comparison with RenderOcc. First, we treat each voxel center as the center of a Gaussian and predict the offset of their properties within the Gaussian. Next, we render the semantics logits and depth and optimize the network with LiDAR rendering supervision. Finally, we introduce a multi-frame calibration module aimed at reducing false positive detections caused by duplicate predictions.
IV-B Gaussian Properties Field.
We treat each voxel center as the initial point for a gaussian distribution, as shown in Figure 3. For a given occupancy feature $V$, in addition to predicting the semantic and opacity attributes, we initialize the Gaussian’s mean and scale, and predict offsets for each property, excluding rotation. In structured voxel spaces, we contend that rotating the Gaussian is unnecessary. Instead, adjusting scale suffices to meet the spatial requirements of the entire voxel space. The regular arrangement of voxels in a uniformly distributed and structured environment already provides sufficient geometric and spatial information, rendering rotational adjustments redundant. Our ablation studies have confirmed this approach. The resulting Gaussian properties $G\in\mathbb{R}^{N\times(7+(L-1))}$, where $N$ denotes the total number of Gaussians. For each Gaussian, we have
@@ IV Method
During the testing phase, we treat Gaussians with opacity greater than a predefined threshold as the final predictions. Specifically
$V(\mu)=\begin{cases}\arg\max\left(C(\mu)\right)&\text{if }o(\mu)\geq\tau\\ \varnothing&\text{if }o(\mu)<\tau\end{cases}$ (8)
where $\tau$ is the opacity threshold for determining whether a Gaussian is considered empty. Since empty voxels are filtered out based on threshold, the actual number of predicted semantic categories is L1L-1, excluding the ’free’ category. Visualization analysis shows that the majority of effective Gaussians closely align with their voxel centers. Thus, we apply the Gaussian semantics directly to the voxels.
IV-B1 Gaussian Rasterizer.
We interpret the semantic logits associated with each Gaussian as colors, and apply alpha-blending to render the semantic logits and depth for each pixel. This approach culminates in the generation of a semantic map $S$ and a depth map $D$.
@@
IV-C Ray Compensation Module.
RenderOcc [pan2024renderocc] introduces Auxiliary Rays to ensure spatial consistency across different viewpoints by aligning labeled rays from multiple frames to a single frame. Nevertheless, it fails to resolve the duplicated predictions along these rays. We introduce the Ray Compensation module, which provides multiple perspectives for each object at the feature level. Specifically as shown in Figure 4. The upper part of figure illustrates the principle of our method. For the current frame $t$, trees that are obscured might be revealed in adjacent frames from the different viewpoints. To obtain features from adjacent time steps, we first assume that the current Occupancy Feature is denoted as $P_{\text{curr}}\in\mathbb{R}^{H_{V}\times W_{V}\times Z_{V}\times C}$. We treat the centers of voxels as homogeneous coordinates $O^{\text{curr\_ego}}_{\text{curr}}\in\mathbb{R}^{H_{V}\times W_{V}\times Z_{V}\times 4}$, representing the voxel centers in the ego-car coordinate system at the current time. The position of the adjacent time step in the current time step’s coordinate system is computed as follows:
@@
IV-C1 Dynamic Object.
It is noteworthy that this approach may still encounter issues, as illustrated in the lower part of Figure 4. For trees obscured in the current frame, they may still be blocked by other moving objects in adjacent frames. In extreme cases, these obstructions might occur at every time step. Therefore, we need to minimize the influence of dynamic objects in adjacent frames. This is achieved by applying a weighting factor, alpha, when calculating the loss for adjacent frames, as detailed in the following subsection.
IV-D Loss Function
To maintain consistency with RenderOcc, we also optimize the Gaussian properties using a class-balanced cross-entropy loss $\mathcal{L}^{\text{seg}}$ for occupancy prediction and the SILog loss $\mathcal{L}^{\text{depth}}$ [eigen2014depth] for depth estimation. For the current frame, we compute the loss similarly to RenderOcc as follows:
$\mathcal{L}_{curr}=\mathcal{L}^{seg}_{curr}+\mathcal{L}^{depth}_{curr}.$ (11)